factor by grouping with complete solution
1. x² - y² - x –y
2. x³ - y³ - x² + 2xy – y²
3. x² + xy – 2y² - x³ + y³

Answer by venugopalramana(3286)

factor by grouping with complete solution
1. x² - y² - x –y
=(X+Y)(X-Y)-1(X+Y)=(X+Y)(X-Y-1)
2. x³ - y³ - x² + 2xy – y²
(X-Y)(X^2+XY+Y^2)-(X-Y)^2
=(X-Y)(X^2+XY+Y^2-X+Y)

3. x² + xy – 2y² - x³ + y³
{X^2+2XY-XY-2Y^2}-(X^3-Y^3)
={X(X+2Y)-Y(X+2Y)}-(X-Y)(X^2+XY+Y^2)
(X-Y)(X+2Y)-(X-Y)(X^2+XY+Y^2)
(X-Y)(X+2Y-X^2-XY-Y^2)